# Relative Lonely Runner spectra

### v1

@@ 1.1. The Lonely Runner Conjecture
The equaility case mentioned at the end of the first paragraph gives $D(\langle(1,2,\ldots,n)\rangle_{\mathbb{R}})=1/2-1/(n+1)\in\mathcal{S}_{1}(n)$, and the difficult part of the Lonely Runner Conjecture is showing that for every $1$-dimensional proper subtorus $T\subseteq(\mathbb{R}/\mathbb{Z})^{n}$, we have $D(T)\leq 1/2-1/(n+1)$.
@@ 1.3. Main result
We make three remarks about the statement of this theorem. First, the finite symmetric difference arises because a few $1$-dimensional subtori of $U$ can be too “small” to see the structure leading to the sets $D(U)+1/\operatorname{Prog}(\alpha_{i},\beta_{i})$. [^1] In our numerical examples (described below), we have not run across any instances where this finite symmetric difference is actually necessary, but we do expect it to appear in general. Second, our argument actually shows the ostensibly slightly stronger statement that there are only finitely many $1$-dimensional proper subtori $T$ of $U$ with $D(T)$ lying outside of the set $\{D(U)\}\cup[\cup_{i=1}^{N}(D(U)+1/\operatorname{Prog}(\alpha_{i},\beta_{i}))]$. Third, the set $D(U)+1/\operatorname{Prog}(\alpha_{i},\beta_{i})$ can be understood as a simple transformation of the set $\mathcal{S}_{1}(2)=1/\operatorname{Prog}(4,6)$.
@@ 1.7. Organization of the paper
We prove Theorem 1.1 in Section 2. In Section 3 we describe some techniques for determining all of the $2$-dimensional subtori with a given $D$-value and carry out this determination for the cases that will figure in our explicit examples. We then prove Theorems 1.3, 1.4, and 1.5 in Sections 4, 5, and 6.

### v2

@@ 1.1. The Lonely Runner Conjecture
The equality case mentioned at the end of the first paragraph gives $D(\langle(1,2,\ldots,n)\rangle_{\mathbb{R}})=1/2-1/(n+1)\in\mathcal{S}_{1}(n)$, and the difficult part of the Lonely Runner Conjecture is showing that for every $1$-dimensional proper subtorus $T\subseteq(\mathbb{R}/\mathbb{Z})^{n}$, we have $D(T)\leq 1/2-1/(n+1)$.
@@ 1.3. Main result
We make three remarks about the statement of this theorem. First, the finite symmetric difference arises because a few $1$-dimensional subtori of $U$ can be too “small” to see the structure leading to the sets $D(U)+1/\operatorname{Prog}(\alpha_{i},\beta_{i})$; below we give an explicit example where this finite symmetric difference is necessary. Second, our argument actually shows the ostensibly slightly stronger statement that there are only finitely many $1$-dimensional proper subtori $T$ of $U$ with $D(T)$ lying outside of the set $\{D(U)\}\cup[\cup_{i=1}^{N}(D(U)+1/\operatorname{Prog}(\alpha_{i},\beta_{i}))]$. Third, the set $D(U)+1/\operatorname{Prog}(\alpha_{i},\beta_{i})$ can be understood as a simple transformation of the set $\mathcal{S}_{1}(2)=1/\operatorname{Prog}(4,6)$.
@@ 1.5. Computing relative spectra
In a related direction, we give a necessary and sufficient geometric condition (Proposition 7.1) for the relative spectrum of a $2$-dimensional subtorus being finite, i.e., having no progressions.
@@ 1.7. Organization of the paper
We prove Theorem 1.1 in Section 2. In Section 3 we describe some techniques for determining all of the $2$-dimensional subtori with a given $D$-value and carry out this determination for the cases that will figure in our explicit examples. We then prove Theorems 1.3, 1.4, and 1.5 in Sections 4, 5, and 6. In Section 7 we establish a criterion for the finiteness of a relative spectrum and give an example of a finite, non-empty relative spectrum.
@@ 6. A new progression in the spectrum $\mathcal{S}_{1}(6)$
7. Finite and infinite relative spectra
In this final section, we make a few brief remarks about when relative spectra are finite. Let $n\geq 3$, and let $U\subseteq(\mathbb{R}/\mathbb{Z})^{n}$ be a $2$-dimensional proper subtorus. Write
$Z(U):=\{x\in U:D(x)=D(U)\}$
for the locus of points where $U$ achieves its $D$-value. Notice that $Z(U)$ is the intersection of $U$ with the hollow cube $\{x\in(\mathbb{R}/\mathbb{Z})^{n}:D(x)=D(U)\}$. We can express $Z(U)$ as the union of the intersections of $U$ with the various facets of this hollow cube. Each such intersection has dimension at most $1$ (otherwise $U$ would be contained in a translate of a coordinate hyperplane), and we conclude that $Z(U)$ consists of a finite union of points and (closed) line segments. Of course these points and the endpoints of these segments all have rational coordinates.
Our criterion for the finiteness of $\mathcal{S}_{1}(U)$ concerns whether or not $Z(U)$ contains non-parallel line segments; notice that the “parallel-ness” of line segments is the same in $(\mathbb{R}/\mathbb{Z})^{n}$ and in $U$ (considered as a $2$-dimensional torus).
Proposition 7.1.
Let $n\geq 3$ be a natural number, and let $U\subseteq(\mathbb{R}/\mathbb{Z})^{n}$ be a $2$-dimensional proper subtorus. The relative spectrum $\mathcal{S}_{1}(U)$ is finite if and only if the set $Z(U)$ contains non-parallel line segments.
Proof.
First, suppose that $Z(U)$ contains two non-parallel line segments. To show that $\mathcal{S}_{1}(U)$ is finite, it suffices to show that all but finitely many $1$-dimensional subtori $T\subseteq U$ intersect the union of these line segments and thus satisfy $D(T)=D(U)$. Fix an identification of $U$ with $(\mathbb{R}/\mathbb{Z})^{2}$, and consider the images $s_{1},s_{2}$ of the two non-parallel line segments. Let $\ell$ be the minimum of the lengths of $s_{1},s_{2}$, and let $L_{1},L_{2}$ denote the lines containing $s_{1},s_{2}$ (respectively). Let $20<\theta\leq\pi/$ 2 be the measure of the angle formed by $L_{1},L_{2}$.
Now, let $T^{\prime}\subseteq(\mathbb{R}/\mathbb{Z})^{2}$ be a $1$-dimensional subtorus. It forms an angle of at least $\theta/2$ with some $L_{i}$. Then the height of $s_{i}$ in the direction perpendicular to $T^{\prime}$ is at least
$\ell\sin(\theta/2).$
The perpendicular distance between consecutive leaves of $T^{\prime}$ is $1/\operatorname{vol}(T^{\prime})$, where $\operatorname{vol}(T^{\prime})$ denotes the volume (i.e., length) of $T^{\prime}$. Thus, if $\operatorname{vol}(T^{\prime})\geq 1/(\ell\sin(\theta/2))$, then $T^{\prime}$ must intersect $s_{i}$, and we conclude that the corresponding $1$-dimensional subtorus $T\subseteq U$ has $D(T)=D(U)$. Since there are only finitely many $1$-dimensional subtori with volume below any given threshold, it follows that all but finitely many $1$-dimensional subtori $T\subseteq U$ have $D(T)=D(U)$, as desired.
Second, suppose that $Z(U)$ does not contain any non-parallel line segments. To show that $\mathcal{S}_{1}(U)$ is infinite, it suffices to show that there are infinitely many $1$-dimensional subtori $T\subseteq U$ that are disjoint from $Z(U)$: Indeed, such $T$’s have $D(T)>D(U)$, and there are infinitely many distinct such values since $D(T)\to D(U)$ as $\operatorname{vol}(T)\to\infty$. By assumption, $Z(U)$ is contained in a finite union of rational parallel lines and hence is contained in a $1$-dimensional subgroup $\Gamma$ of $U$.
We claim that $U$ contains at least two non-parallel $1$-dimensional subtori that are contained in the coordinate hyperplanes of $(\mathbb{R}/\mathbb{Z})^{n}$. Indeed, for each $1\leq i\leq n$, the intersection $U\cap\{x_{i}=0\}$ is a $1$-dimensional subgroup, and the identity components of these intersections cannot all be identical because then this common identity component would consist of only the point 00. Thus, there is a $1$-dimensional subtorus $T^{*}$ of $U$ that is contained in a coordinate hyperplane of $(\mathbb{R}/\mathbb{Z})^{n}$ and is not parallel to $\Gamma$. Fix an identification of $U$ with $(\mathbb{R}/\mathbb{Z})^{2}$ such that $T^{*}$ is identified with the $x_{2}$-axis; let $\Gamma^{\prime}$ denote the image of $\Gamma$ under this identification. Since $D(T^{*})=1/2>D(U)$, we know that $Z(U)$ is disjoint from $T^{*}$, and it follows that the image of $Z(U)$ is contained in $\Gamma^{\prime}\setminus(\{0\}\times(\mathbb{R}/\mathbb{Z}))$. Hence it suffices to find infinitely many $1$-dimensional subtori $T\subseteq(\mathbb{R}/\mathbb{Z})^{2}$ whose intersections with $\Gamma^{\prime}$ are contained in $\{0\}\times(\mathbb{R}/\mathbb{Z})$.
The choice of $T^{*}$ guarantees that $\Gamma^{\prime}$ has some finite slope $b/a$, where $a,b$ are coprime integers with $a>0$. Let $q\in\mathbb{N}$ be such that the leaves of $\Gamma^{\prime}$ are spaced out $1/q$ apart vertically. Now, for each $s\in\mathbb{N}$, consider the $1$-dimensional subtorus
$T_{s}:=\langle(a,b+a/(qs))\rangle_{\mathbb{R}}.$
As we travel along $T_{s}$ starting at $(0,0)$, the first intersection with $\Gamma^{\prime}$ occurs at the point
$(s/a)(a,b+1/(qs))=(s,sb/a+1/q)\in\{0\}\times(\mathbb{R}/\mathbb{Z}),$
so in fact $T_{s}\cap\Gamma^{\prime}\subseteq\{0\}\times(\mathbb{R}/\mathbb{Z})$, as desired. ∎
We remark that [9] established $\mathcal{S}_{1}(n-1)\subset\operatorname{acc}\mathcal{S}_{1}(n)$ by showing that $\mathcal{S}_{1}(U^{\prime}\times(\mathbb{R}/\mathbb{Z}))$ is infinite for every proper $1$-dimensional subtorus $U^{\prime}\subseteq(\mathbb{R}/\mathbb{Z})^{n-1}$; the construction in the second part of the proof of Proposition 7.1 is a generalization of this argument.
It is not a priori obvious that the geometric condition for finiteness in Proposition 7.1 is ever satisfied. In fact, one can show by ad-hoc arguments that it is never satisfied for $n\leq 6$. For $n=7$, however, there are examples like the following. Consider the $2$-dimensional proper subtorus
$U^{8}:=\langle(1,2,3,2,0,0,0),(0,0,0,2,1,2,3)\rangle_{\mathbb{R}}.$
A tedious but routine calculation gives that $D(U^{8})=3/10$, and that $Z(U^{8})$ consists of the points $a(1,2,3,2,0,0,0)+b(0,0,0,2,1,2,3)$ where $(a,b)$ either lies on one of the eight line segments
$\{2/5,3/5\}\times([1/5,4/15]\cup[11/15,4/5]),\quad([1/5,4/15]\cup[11/15,4/5])\times\{2/5,3/5\}.$
or is one of the four points $(1/5,1/5),(2/5,2/5),(3/5,3/5),(4/5,4/5)$. It follows from Proposition 7.1 that $\mathcal{S}_{1}(U^{8})$ is finite, i.e., $\mathcal{S}_{1}(U^{8})$ does not contain any progressions. Using a computer to calculate $D(T)$ for low-volume proper $1$-dimensional subtori $T\subseteq U^{8}$ reveals that $\mathcal{S}_{1}(U^{8})$ contains (at least) the values $7/22,11/34,17/54,23/74$ in addition to $3/10$. Hence, for $\mathcal{S}_{1}(U^{8})$, the finite symmetric difference in Theorem 1.1 is necessary.